Do you know the pattern or formula or this equation?
(x-1)(x-2)=0
Thanks!:)

Answer by jim_thompson5910(35256)

(x-1)(x-2)=0

x-1=0 or x-2=0 ... Use the zero product property

x-1+1=0+1 or x-2+2=0+2

x+0=1 or x+0=2

x=1 or x=2

So the two solutions are x=1 or x=2